If 60% people in city like cricket, 30% like football and the remaining like other games, then what per cent of the people like other games? If the total number of people are 50 lakh, find the exact number who like each type of game. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Percentage of people who like other games= (100 − 60 − 30)%
= (100 − 90)% = 10 %
Total number of people= 50 lakh
Number of people who like cricket =60/100 × 50 = 30 lakh
Number of people who like football =30/100 × 50 = 15 lakh
Number of people who like other games=10/100 × 50= 5 lakh

Thanks